OpenStudy (ssprague_77):

A = 3(B + C) solve for B

OpenStudy (ssprague_77):

This is not a test question, the last one was. but i finished it.

563blackghost (563blackghost):

We need to isolate B.... \(\huge{A=3(B+C)}\) We would first divide by 3 to get it to the other side (what we do on one side we do to the other). \(\huge{\frac{A}{3}=\frac{3}{3}(B+C)}\)

563blackghost (563blackghost):

This would simplify to.... \(\huge{\frac{A}{3}=(B+C)}\) What would we do next?

OpenStudy (ssprague_77):

We would then combine like terms? I'm kinda bad in Algebra 1... sorry

563blackghost (563blackghost):

Not quite. In our equation we do not have any like terms...we have three different variables and a real number so no like terms....since we need to isolate B we would subtract C what we do on one side we do to the other.... \(\huge{-C+\frac{A}{3}=B(C-C)} \rightarrow \color{red}{-C+\frac{A}{3}=B}\)

OpenStudy (ssprague_77):

oh okay.... I think I get it now. Thanks!

563blackghost (563blackghost):

np ;)